# Differentially private projection-depth-based medians

### v2

@@ 3.2 Verifying Condition 1 and example applications
Suppose there exists $40<q^{\prime}<q<1/$ 4 such that both $6<n(q-q^{\prime})$ and Condition 4 holds for $(q^{\prime},\nu)$. Then the median absolute deviation satisfy Condition 1 with $a_{1}\lesssim n\exp(d)$, $a_{2}=3M/24$ and $\kappa_{n}=\exp({-nr^{2}M^{2}+d\log(cn/d)}).$
The proofs are deferred to Appendix C, and follow mainly from Lemma C.2. Theorem 3.4 and Theorem 3.5 can be interpreted as “provided the (univariate) projected densities are bounded below when computed at the inner quantiles, then the trimmed estimators and the median absolute deviation have small projected sensitivity with high probability.” Note that we can prove an analogue of Theorem 3.5 with the interquartile range in place of the median absolute deviation (see Lemma C.5). To illustrate the methodology, we begin with a simple example, taking $\nu$ to be Gaussian.
@@
$\zeta_{n}(t)\lesssim\exp\left(-cnt^{2}/\beta_{d}+d\right),$
@@
$\zeta_{n}(t)=\exp\left(-c^{\prime}nt^{2}/d+d\log(cn/d)\right)+\exp\left(-c^{\prime}n/d^{2}+d\log(cn/d)\right),$
@@
$\mathbb{E}\lambda(\hat{\nu},\epsilon,\delta)$ $\gtrsim 1-e^{\epsilon}\delta-\exp\left(-\frac{n\eta}{d^{1/2}\tau}+d+\log n\right)-\exp\left(-\frac{n}{d^{2}}+d\log(n/d)\right).$
Thus, compared to the Gaussian case, the lower bound on $\delta$ is larger. Furthermore, we require $\eta\gtrsim d^{3/2}/n\vee d^{1/2}\log n/n$ to ensure that $\mathbb{E}\lambda(\hat{\nu},\epsilon,\delta)\gtrsim n^{-L}$, a factor of $d^{1/2}$ larger than in the Gaussian case. We can now apply Theorem 3.2. Taking $\eta\propto d^{3/2}/n$, for $n^{-L}<\gamma<1$ with probability $1-\gamma$, it holds that
$\lVert\tilde{\theta}-\theta\rVert\lesssim\frac{d^{3/2}\left[\log(1/\gamma)\vee d\log d\right]}{n\epsilon}\vee\sqrt{\frac{d^{2}\log(1/\gamma)\vee d^{3}\log(n/d)}{n}}.$
It is immediate that the scaling in the dimension of the deviations bound is worse than that of the Gaussian case (Example 3.1). This is expected, given that heavy-tailed estimation is a more challenging problem. We refer to this as “outlier error amplification”. A similar phenomenon, in the form of sample complexity bounds, was observed by Ramsay et al., [21], for private medians based on other depth functions. However, it is important to note that the obtained bound is only an upper bound on the deviations of the estimator. Indeed, to the best of our knowledge, the lower bound on the sample complexity of differentially private estimation of the location parameter from a product of Cauchy marginals is unknown. It is straightforward to show that this is at least as hard as differentially private Gaussian mean estimation, see Lemma F.1. However, there is gap between the upper bounds derived here and in Ramsay et al., [21] and the lower bound for Gaussian mean estimation. We conjecture that a tighter lower bound exists, which matches the upper bound of Ramsay et al., [21]. This is supported by the numerical experiments we completed, see Section 4, where we observed the “outlier error amplification” effect.
@@ 5 General results on propose-test-release mechanisms
In this section, we elaborate on the notion of safety [6, 17] in the general case of PTR mechanisms. We show that safety is both sufficient and necessary for privacy of a PTR mechanism. We also present general results for analyzing a PTR mechanism based on the exponential mechanism. Let $Ber(p)$ denote the Bernoulli measure with parameter $0\leq p\leq 1$.
@@
A test mechanism is any mechanism, $Q$, such that for all $\hat{\nu}\in\widehat{\mathcal{M}}_{1}(n,d)$, $Q_{\hat{\nu}}=Ber(\lambda(\hat{\nu}))$ for some $0\leq\lambda(\hat{\nu})\leq 1$.
Every test mechanism is characterized by its associated $\lambda$. We will refer to a test mechanism by its associated $\lambda$ function.
@@
Given a mechanism $Q$ mapping to $\mathcal{M}_{1}(\Theta)$ and a test mechanism $\lambda$, the PTR mechanism based on $(Q,\lambda)$ is given by the map
@@
The probability of any event under the PTR mechanism will be the probability that the test is passed multiplied by the probability of the event under the original mechanism, plus the probability of the test failing if $\perp$ is an element of the event. We now introduce the notions of safety, safety margin, and the exponential mechanism in the general case.
@@
Given $\epsilon,\delta>0$, the pair $(Q,\hat{\nu})$ is $(\epsilon,\delta)$-safe if, for all $\tilde{\nu}\in\widetilde{\mathcal{M}}(\hat{\nu},1)$ and $Q$-measurable sets $B$, it holds that
$Q_{\hat{\nu}}(B;\epsilon,\delta)\leq e^{\epsilon}Q_{\tilde{\nu}}(B;\epsilon,\delta)+\delta\qquad\text{and}\qquad Q_{\tilde{\nu}}(B;\epsilon,\delta)\leq e^{\epsilon}Q_{\hat{\nu}}(B;\epsilon,\delta)+\delta.$
@@
We include $\epsilon$ and $\delta$ in the notation for the mechanism $Q$ to emphasize that $Q$ typically depends on $\epsilon$ and $\delta$. Safety is linked to the privacy of a given PTR mechanism. Colloquially, if $\lambda(\hat{\nu})$ is large, then the dataset must satisfy some level of safety. In addition, sufficient conditions for differential privacy of a PTR mechanism are that the test mechanism is differentially private and that it satisfies $\lambda(\hat{\nu})$ is very small if $\hat{\nu}$ is unsafe.
@@
When $\epsilon$ and $\delta$ are clear from the context, we omit them, denoting the safety margin as $\operatorname{SM}(\hat{\nu},Q).$ If $\lambda(\hat{\nu})$ is small whenever $\hat{\nu}$ is not in the safe set of $Q$ and the test mechanism is differentially private, then a PTR mechanism will be differentially private.
@@
Given $\epsilon_{1},\epsilon_{2},\delta_{1},\delta_{2},\delta_{3}>0$, let $P$ be the PTR mechanism based on $Q(\cdot;\cdot,\epsilon_{2},\delta_{2})$ and $\lambda(\cdot,\epsilon_{1},\delta_{1})$. If $\lambda$ is $(\epsilon_{1},\delta_{1})$-differentially private and is such that for any $\hat{\nu}\notin\mathbf{S}(Q,\epsilon_{2},\delta_{2})$, $\lambda(\hat{\nu},\epsilon_{1},\delta_{1})<\delta_{3}$, then $P$ is $(\epsilon_{1}+\epsilon_{2},\delta_{1}+\delta_{2}\vee\delta_{3})$-differentially private.
@@
Suppose that a PTR mechanism $P$ based on $Q$ and $\lambda$ is $(\epsilon,\delta)$-differentially private. Then
@@
    For any $\hat{\nu}$ such that $\lambda(\hat{\nu})>1/k$, the pair $(Q(;\hat{\nu},\epsilon),\hat{\nu})$ is $(2\epsilon,k\delta e^{\epsilon})$-safe.
@@
For a cost function $\phi(\cdot;\hat{\nu})$, let $A_{\tau,\hat{\nu}}=\{\phi(x,\hat{\nu})\leq\tau\}$. The exponential mechanism $\operatorname{EM}_{\hat{\nu},\phi}(\cdot;\tau,\eta,\epsilon)$ is the measure whose density satisfies
@@
Now, the critical components for accuracy are that $\Pr\left(\hat{\nu}\in\mathbf{S}(Q,\epsilon,\delta)\right)$ is large and that $\lambda(\hat{\nu})$ is large when $\hat{\nu}\in\mathbf{S}(Q,\epsilon,\delta)$. We formalize this for the well-known exponential mechanism as follows. Let $G$ be the CDF of a standard Laplace RV.
@@ Appendix B Proofs from Section 3
    $\Pr\Big{(}\inf_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ \tilde{\nu}\in\{\hat{\nu}_{u}\}\cup\widetilde{\mathcal{M}}_{1}(\hat{\nu}_{u},k)\end{subarray}}\sigma(\tilde{\nu})\leq\frac{c_{1}}{2}\Big{)}\lesssim a_{1}e^{-a_{2}nc_{1}/4}+\kappa_{n}+\zeta_{n}(\frac{c_{1}}{4}).$ (B.1)
@@
    $\Pr\left(\hat{\alpha}_{\mu}>4c_{3}\right)\lesssim\zeta_{n}(c_{3}).$ (B.2)
@@
    $\Pr\left(\sup_{u\in\mathbb{S}^{d-1}}\sigma(\hat{\nu}_{u})\geq 2c_{2}\right)\lesssim\zeta_{n}(c_{2}).$ (B.3)
@@
    $\Pr\left(\inf_{u\in\mathbb{S}^{d-1}}\sigma(\hat{\nu}_{u})\leq c_{1}/2\right)\lesssim\zeta_{n}(\frac{c_{1}}{2}).$ (B.4)
@@
$\lesssim a_{1}e^{-a_{2}nc_{1}/4}+\kappa_{n}+\zeta_{n}(\frac{c_{1}}{4}).$
@@
$\Pr\left(\hat{\alpha}_{\mu}\geq 4c_{3}\right)$ $\leq\Pr\left(2\left(\sup_{u\in\mathbb{S}^{d-1}}|\mu_{u}-\hat{\mu}_{u}|+c_{3}\right)>4c_{3}\right)\lesssim\zeta_{n}(c_{3}).$
@@
$\lesssim\zeta_{n}(\frac{c_{1}}{2}),$
@@
First, note that $\mu$ and $\sigma$ may satisfy Conditions 1 and 2 for different parameters, e.g., different sequences $\zeta_{n}$. However, we can define a set of joint parameters $\zeta_{n},\kappa_{n},\ldots$ for which the bounds in Conditions 1 and 2 hold for both $\mu$ and $\sigma$, which we will use throughout the proof.
@@
$\qquad\lesssim a_{1}\exp\left(-a_{2}\frac{c_{1}n}{4}\left[1\wedge\frac{\eta}{\tau\vee 1}\right]\right)+\kappa_{n}+\zeta_{n}(\frac{c_{1}}{4}).$
@@
$\Pr\left(E_{5}^{c}\right)\lesssim a_{1}e^{-a_{2}nc_{1}/4}+\kappa_{n}+\zeta_{n}(\frac{c_{1}}{4}\wedge c_{3}).$
@@
$\mathbb{E}Z$ $\geq 1-\Pr\left(E_{1}\right)-\Pr\Big{(}\sup_{\begin{subarray}{c}\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},3)\\ x\in A_{\tau+\eta,\hat{\nu}}\cup A_{\tau+\eta,\tilde{\nu}}\end{subarray}}|O(x,\hat{\nu})-O(x,\tilde{\nu})|>\eta\Big{)}-\delta$
$\gtrsim 1-\Pr\left(E_{5}^{c}\right)-\left(a_{1}\exp\left(-a_{2}\frac{c_{1}n}{4}\left[1\wedge\frac{\eta}{\tau\vee 1}\right]\right)+\kappa_{n}+\zeta_{n}(\frac{c_{1}}{4})\right)-\delta$
$\gtrsim 1-(a_{1}e^{-a_{2}nc_{1}/4}+\kappa_{n}+\zeta_{n}(\frac{c_{1}}{4}\wedge c_{3}))$
$\qquad-\left(a_{1}\exp\left(-a_{2}\frac{c_{1}n}{4}\left[1\wedge\frac{\eta}{\tau\vee 1}\right]\right)+\kappa_{n}+\zeta_{n}(\frac{c_{1}}{4})\right)-\delta$
$\gtrsim 1-\exp\left(-a_{2}\frac{c_{1}n}{4}\left[1\wedge\frac{\eta}{\tau\vee 1}\right]\right)-\kappa_{n}-\zeta_{n}(\frac{c_{1}}{4})-\delta\qed.$
@@
$\Pr\left(\lVert\tilde{\theta}-\theta\rVert\geq t\big{|}\tilde{\theta}\neq\perp\right)\\ \lesssim\inf_{\begin{subarray}{c}0<s<c_{1}\tau/2-c_{3}\\ 0<r<c_{1}\tau/2-s-c_{3}\end{subarray}}\Bigg{[}\exp\left(\epsilon\frac{(c_{1}/2+2c_{2})(c_{3}+s)}{4\eta c_{1}c_{2}}-\frac{\epsilon t}{8\eta c_{2}}+\frac{\epsilon r}{2c_{1}\eta}+\log\frac{(2c_{2}\tau+s+c_{3})^{d}-t^{d}}{r^{d}}\right)\\ +\zeta_{n}\left(s\wedge c_{3}\wedge c_{1}/2\right)\Bigg{]}.$
@@
Condition 2 directly gives that $\Pr\left(E_{6,s}^{c}\right)\lesssim\zeta_{n}(s)$. Using this, the identities established in (B.2), (B.3), and (B.4), and noting that $c_{1}\leq c_{2}$, yields that for all $s\geq 0$, we have
$\Pr\left((E_{7}\cap E_{6,s})^{c}\right)\lesssim\zeta_{n}(s\wedge c_{3}\wedge c_{1}/2).$ (B.8)
@@
$\Pr\left(\lVert\tilde{\theta}-\theta\rVert\geq t\big{|}\tilde{\theta}\neq\perp\right)\\ \lesssim\inf_{\begin{subarray}{c}0<s<c_{1}\tau/2-c_{3}\\ 0<r<c_{1}\tau/2-s-c_{3}\end{subarray}}\Bigg{[}\exp\left(\frac{\epsilon(c_{1}/2+2c_{2})(c_{3}+s)}{4\eta c_{1}c_{2}}-\frac{\epsilon t}{8\eta c_{2}}+\frac{\epsilon r}{2c_{1}\eta}+\log\frac{(2c_{2}\tau+s+c_{3})^{d}-t^{d}}{r^{d}}\right)\\ +\zeta_{n}\left(s\wedge c_{3}\wedge c_{1}/2\right)\Bigg{]}.$ (B.12)
@@
Next we bound $\zeta_{n}(s\wedge c_{3}\wedge c_{1}/2)\leq\gamma/3$. By assumption, $\gamma>3\zeta_{n}(c_{1}/2\wedge c_{3})$, and so we consider only $\zeta_{n}(s)$ for $s<c_{1}/2\wedge c_{3}$. To this end, using the fact that $\zeta_{n}$ are non-increasing, we have that
$\zeta_{n}(s)\leq\gamma/3\implies t\gtrsim\frac{c_{2}}{c_{1}}\zeta_{n}^{-1}(\gamma/3).$
@@
Before proving Theorems 3.4 and 3.5, we first prove two results concerning the projected quantiles and spacings. Let $\lambda_{t,q_{1},q_{2}}=\inf_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ z\in(\xi_{q_{1},u}-t,\xi_{q_{2},u}+t)\end{subarray}}f_{\nu_{u}}(z)$ and $\hat{\xi}_{q,u}=F^{-1}_{\hat{\nu}_{u}}(q)$.
@@
We now consider the event $E_{9}$. Recall that the set $\mathcal{N}_{\epsilon}$ is an $\epsilon$-net on $\mathbb{S}^{d-1}$ if for all $u\in\mathbb{S}^{d-1}$, there exists $v\in\mathcal{N}_{\epsilon}$ such that $\lVert u-v\rVert\leq\epsilon$. Letting $\mathcal{N}_{\epsilon}$ be an arbitrary $\epsilon$-net on $\mathbb{S}^{d-1}$, it is straightforward (see, e.g., [26, Lemma 4.4.1]) to show that for any $i,j\in[n]$,
@@
Furthermore, there exists $\mathcal{N}_{1/4}$ such that $|\mathcal{N}_{1/4}|\leq 9^{d}$ [26, see Corollary 4.2.13]. Applying these two facts, in conjunction with two applications of subadditivity of probability measures, yields

### v3

@@ 3.2 Verifying Condition 1 and example applications
Suppose there exists $40<q^{\prime}<q<1/$ 4 such that both $6<n(q-q^{\prime})$ and Condition 4 holds for $(q^{\prime},\nu)$. Then the median absolute deviation satisfies Condition 1 with $a_{1}\lesssim n\exp(d)$, $a_{2}=3M/24$ and $\kappa_{n}=\exp({-nr^{2}M^{2}+d\log(cn/d)}).$
The proofs are deferred to Appendix C, and follow mainly from Lemma C.2. Theorem 3.5 and Theorem 3.6 can be interpreted as stating that, provided the (univariate) projected densities are bounded below when computed at the inner quantiles, the trimmed estimators and the median absolute deviation have small projected sensitivity with high probability. Note that we can prove an analogue of Theorem 3.6 with the interquartile range in place of the median absolute deviation (see Lemma C.5). To illustrate the methodology, we begin with a simple example, taking $\nu$ to be Gaussian.
@@
$\zeta_{n,\mu}(t)=\zeta_{n,\sigma}(t)=\exp\left(-cnt^{2}/\beta_{d}+d\right),$
@@
$\zeta_{n,\mu}(t)=\exp\left(-2nt^{2}/d+d\log(cn/d)\right),$
and
$\zeta_{n,\sigma}(t)=\exp\left(-c^{\prime}nt^{2}/d+d\log(cn/d)\right)+\exp\left(-c^{\prime}n/d^{2}+d\log(cn/d)\right),$
@@
$\mathbb{E}Z$ $\geq 1-(e^{\epsilon}\delta-\exp\left(-\frac{n\eta}{d^{1/2}\tau}+d+\log n\right)-\exp\left(-\frac{n}{d^{2}}+d\log(n/d)\right).$
Thus, compared to the Gaussian case, the lower bound on $\delta$ is larger. Furthermore, we require $\eta\gtrsim d^{3/2}/n\vee d^{1/2}\log n/n$ ensuring that $\mathbb{E}Z\gtrsim n^{-L}$, a factor of $d^{1/2}$ larger than in the Gaussian case. We can now apply Theorem 3.2. Taking $\eta\propto d^{3/2}/n$, for $n^{-L}<\gamma<1$ with probability $1-\gamma$, it holds that
$\lVert\tilde{\theta}-\theta\rVert\lesssim\frac{d^{3/2}\left[\log(1/\gamma)\vee d\log d\right]}{n\epsilon}\vee\sqrt{\frac{d\log(1/\gamma)\vee d^{2}\log(n/d)}{n}}.$
It is immediate that the scaling in the dimension of the deviations bound is worse than that of the Gaussian case (Example 3.1). This is expected, given that heavy-tailed estimation is a more challenging problem. We refer to this as “outlier error amplification”. A similar phenomenon, in the form of sample complexity bounds, was observed by Ramsay et al., [24], for private medians based on other depth functions. However, it is important to note that the obtained bound is only an upper bound on the deviations of the estimator. Indeed, to the best of our knowledge, the lower bound on the sample complexity of differentially private estimation of the location parameter from a product of Cauchy marginals is unknown. It is straightforward to show that this is at least as hard as differentially private Gaussian mean estimation, see Lemma F.1. However, there is a gap between the upper bounds derived here and in Ramsay et al., [24] and the lower bound for Gaussian mean estimation. We conjecture that a tighter lower bound exists, which matches the upper bound of Ramsay et al., [24]. This is supported by the numerical experiments we completed, see Section 4, where we observed the “outlier error amplification” effect.
@@ 5 General results on propose-test-release mechanisms
In this section, we elaborate on the notion of safety [6, 19] in the general case of PTR mechanisms. We show that safety is both sufficient and necessary for privacy of a PTR mechanism. We also present general results for analyzing a PTR mechanism based on the exponential mechanism. We begin by introducing a general notion of a test mechanism.
@@
A test mechanism is any mechanism, $Q$, such that for all $\hat{\nu}\in\widehat{\mathcal{M}}_{1}(n,d)$, $Q_{\hat{\nu}}$ is a Bernoulli measure with probability of success $\lambda(\hat{\nu})$, for some $0\leq\lambda(\hat{\nu})\leq 1$.
Definition 5.1 simply says that $Q$ maps an empirical measure $\hat{\nu}$ to a Bernoulli measure with probability of success $\lambda(\hat{\nu})$, for some $0\leq\lambda(\hat{\nu})\leq 1$. Note that every test mechanism is characterized by its associated $\lambda$ function.
@@
Given a mechanism $Q$ whose range is $\mathcal{M}_{1}(\Theta)$ and a test mechanism characterized by $\lambda$, the PTR mechanism based on $(Q,\lambda)$ is given by the map
@@
The probability of any event under the PTR mechanism $P$ applied to $\hat{\nu}$ will be the probability that the test is passed, multiplied by the probability of the event under the mechanism $Q$ applied to $\hat{\nu}$, plus the probability of the test failing if $\perp$ is an element of the event. We now introduce generalized definitions of safety and the safety margin, which apply to arbitrary mechanisms. Subsequently, we will introduce a generalized definition of the exponential mechanism, which applies to an arbitrary cost function $\phi$.
@@
Given $\epsilon,\delta>0$, the pair $(Q,\hat{\nu})$ is $(\epsilon,\delta)$-safe if, for all $\tilde{\nu}\in\widetilde{\mathcal{M}}(\hat{\nu},1)$ and measurable sets $B$, it holds that
$Q_{\hat{\nu}}(B)\leq e^{\epsilon}Q_{\tilde{\nu}}(B)+\delta\qquad\text{and}\qquad Q_{\tilde{\nu}}(B)\leq e^{\epsilon}Q_{\hat{\nu}}(B)+\delta.$
@@
Observe that Definition 5.3 is a generalized version of Definition 3.1. Definition 3.1 is a special case of Definition 5.3, with $Q=\operatorname{EM}$. Note that $Q$ is allowed to depend on $\epsilon$ and $\delta$. Safety is linked to the privacy of a given PTR mechanism. Colloquially, if $\lambda(\hat{\nu})$ is large, then the dataset must satisfy some level of safety. In addition, sufficient conditions for differential privacy of a PTR mechanism are that the test mechanism is differentially private and that $\lambda(\hat{\nu})$ is very small if $\hat{\nu}$ is unsafe.
@@
Observe that Definition 5.4 is a generalized version of Definition 3.2. That is, Definition 3.2 is a special case of Definition 5.4 with $Q$ always set to $\operatorname{EM}$. When $\epsilon$ and $\delta$ are clear from the context, we omit them, denoting the safety margin as $\operatorname{SM}(\hat{\nu},Q).$ If $\lambda(\hat{\nu})$ is small whenever $\hat{\nu}$ is not in the safe set of $Q$ and the test mechanism is differentially private, then a PTR mechanism will be differentially private.
@@
Given $\epsilon_{1},\epsilon_{2},\delta_{1},\delta_{2},\delta_{3}>0$, let $P$ be the PTR mechanism based on $Q$ and $\lambda$. If $\lambda$ is $(\epsilon_{1},\delta_{1})$-differentially private and is such that for any $\hat{\nu}\notin\mathbf{S}(Q,\epsilon_{2},\delta_{2})$, $\lambda(\hat{\nu},\epsilon_{1},\delta_{1})<\delta_{3}$, then $P$ is $(\epsilon_{1}+\epsilon_{2},\delta_{1}+\delta_{2}\vee\delta_{3})$-differentially private.
@@
Suppose that a PTR mechanism $P$ based on $Q$ and $\lambda$ is $(\epsilon,\delta)$-differentially private. Then the follow statements hold:
@@
    For any $\hat{\nu}$ such that $\lambda(\hat{\nu})>1/k$, the pair $(Q,\hat{\nu})$ is $(2\epsilon,k\delta e^{\epsilon})$-safe.
@@
We now consider PTR mechanisms based on general exponential mechanisms. For a general cost function $\phi(\cdot;\hat{\nu})$, let $A_{\tau,\hat{\nu}}=\{\phi(x,\hat{\nu})\leq\tau\}$. Then, we can define the exponential mechanism $\operatorname{EM}_{\hat{\nu},\phi}(\cdot;\tau,\eta,\epsilon)$ as the measure whose density satisfies
@@
Now, the critical components for accuracy are that $\Pr\left(\hat{\nu}\in\mathbf{S}(Q,\epsilon,\delta)\right)$ is large and that $\lambda(\hat{\nu})$ is large when $\hat{\nu}\in\mathbf{S}(Q,\epsilon,\delta)$. We formalize this as follows. Let $G$ be the CDF of a standard Laplace RV.
@@ Appendix B Proofs from Section 3
    $\Pr\Big{(}\inf_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ \tilde{\nu}\in\{\hat{\nu}_{u}\}\cup\widetilde{\mathcal{M}}_{1}(\hat{\nu}_{u},k)\end{subarray}}\sigma(\tilde{\nu})\leq\frac{c_{1}}{2}\Big{)}\lesssim a_{1}e^{-a_{2}nc_{1}/4}+\kappa_{n}+\zeta_{n,\sigma}(\frac{c_{1}}{4}).$ (B.1)
@@
    $\Pr\left(\hat{\alpha}_{\mu}>4c_{3}\right)\lesssim\zeta_{n,\mu}(c_{3}).$ (B.2)
@@
    $\Pr\left(\sup_{u\in\mathbb{S}^{d-1}}\sigma(\hat{\nu}_{u})\geq 2c_{2}\right)\lesssim\zeta_{n,\sigma}(c_{2}).$ (B.3)
@@
    $\Pr\left(\inf_{u\in\mathbb{S}^{d-1}}\sigma(\hat{\nu}_{u})\leq c_{1}/2\right)\lesssim\zeta_{n,\sigma}(\frac{c_{1}}{2}).$ (B.4)
@@
$\lesssim a_{1}e^{-a_{2}nc_{1}/4}+\kappa_{n}+\zeta_{n,\sigma}(\frac{c_{1}}{4}).$
@@
$\Pr\left(\hat{\alpha}_{\mu}\geq 4c_{3}\right)$ $\leq\Pr\left(2\left(\sup_{u\in\mathbb{S}^{d-1}}|\mu_{u}-\hat{\mu}_{u}|+c_{3}\right)>4c_{3}\right)\lesssim\zeta_{n,\mu}(c_{3}).$
@@
$\lesssim\zeta_{n,\sigma}(\frac{c_{1}}{2}),$
@@
First, note that $\mu$ and $\sigma$ may satisfy Conditions 1 for different parameters, e.g., different sequences $\kappa_{n}$. However, we can define a set of joint parameters $\kappa_{n},a_{1},a_{2}$ for which the bound in Condition 1 holds for both $\mu$ and $\sigma$, which we will use throughout the proof.
@@
$\qquad\lesssim a_{1}\exp\left(-a_{2}\frac{c_{1}n}{4}\left[1\wedge\frac{\eta}{\tau\vee 1}\right]\right)+\kappa_{n}+\zeta_{n,\sigma}(\frac{c_{1}}{4}).$
@@
$\Pr\left(E_{5}^{c}\right)\lesssim a_{1}e^{-a_{2}nc_{1}/4}+\kappa_{n}+\zeta_{n,\sigma}(\frac{c_{1}}{4})\wedge\zeta_{n,\mu}(c_{3}).$
@@
$1-\mathbb{E}Z$ $\leq\Pr\left(E_{1}\right)+\Pr\Big{(}\sup_{\begin{subarray}{c}\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},3)\\ x\in A_{\tau+\eta,\hat{\nu}}\cup A_{\tau+\eta,\tilde{\nu}}\end{subarray}}|O(x,\hat{\nu})-O(x,\tilde{\nu})|>\eta\Big{)}+e^{\epsilon}\delta$
$\lesssim\Pr\left(E_{5}^{c}\right)+a_{1}\exp\left(-a_{2}\frac{c_{1}n}{4}\left[1\wedge\frac{\eta}{\tau\vee 1}\right]\right)+\kappa_{n}+\zeta_{n,\sigma}(\frac{c_{1}}{4})+e^{\epsilon}\delta$
$\lesssim a_{1}e^{-a_{2}nc_{1}/4}+\kappa_{n}+\zeta_{n,\sigma}(\frac{c_{1}}{4})\wedge\zeta_{n,\mu}(c_{3})$
$\qquad+a_{1}\exp\left(-a_{2}\frac{c_{1}n}{4}\left[1\wedge\frac{\eta}{\tau\vee 1}\right]\right)+\kappa_{n}+\zeta_{n,\sigma}(\frac{c_{1}}{4})+e^{\epsilon}\delta$
$\lesssim\exp\left(-a_{2}\frac{c_{1}n}{4}\left[1\wedge\frac{\eta}{\tau\vee 1}\right]\right)+\kappa_{n}+\zeta_{n,\sigma}(\frac{c_{1}}{4})+e^{\epsilon}\delta\qed.$
@@
$\Pr\left(\lVert\tilde{\theta}-\theta\rVert\geq t\big{|}\tilde{\theta}\neq\perp\right)\\ \lesssim\inf_{\begin{subarray}{c}0<s<c_{1}\tau/2-c_{3}\\ 0<r<c_{1}\tau/2-s-c_{3}\end{subarray}}\Bigg{[}\exp\left(\epsilon\frac{(c_{1}/2+2c_{2})(c_{3}+s)}{4\eta c_{1}c_{2}}-\frac{\epsilon t}{8\eta c_{2}}+\frac{\epsilon r}{2c_{1}\eta}+\log\frac{(2c_{2}\tau+s+c_{3})^{d}-t^{d}}{r^{d}}\right)\\ +\zeta_{n,\mu}\left(s\wedge c_{3}\right)+\zeta_{n,\sigma}\left(c_{1}/2\right)\Bigg{]}.$
@@
Condition 2 directly gives that $\Pr\left(E_{6,s}^{c}\right)\lesssim\zeta_{n,\mu}(s)$. Using this, the identities established in (B.2), (B.3), and (B.4), and noting that $c_{1}\leq c_{2}$, yields that for all $s\geq 0$, we have
$\Pr\left((E_{7}\cap E_{6,s})^{c}\right)\lesssim\zeta_{n,\mu}(s\wedge c_{3})+\zeta_{n,\mu}(c_{1}/2).$ (B.8)
@@
$\Pr\left(\lVert\tilde{\theta}-\theta\rVert\geq t\big{|}\tilde{\theta}\neq\perp\right)\\ \lesssim\inf_{\begin{subarray}{c}0<s<c_{1}\tau/2-c_{3}\\ 0<r<c_{1}\tau/2-s-c_{3}\end{subarray}}\Bigg{[}\exp\left(\frac{\epsilon(c_{1}/2+2c_{2})(c_{3}+s)}{4\eta c_{1}c_{2}}-\frac{\epsilon t}{8\eta c_{2}}+\frac{\epsilon r}{2c_{1}\eta}+\log\frac{(2c_{2}\tau+s+c_{3})^{d}-t^{d}}{r^{d}}\right)\\ +\zeta_{n,\mu}\left(s\wedge c_{3}\right)+\zeta_{n,\sigma}\left(c_{1}/2\right)\Bigg{]}.$ (B.12)
@@
Next we bound $\zeta_{n,\mu}(s\wedge c_{3})+\zeta_{n,\sigma}(c_{1}/2)\leq\gamma/3$. By assumption, $\gamma>3(zeta_{n,\sigma}(\frac{c_{1}}{2})\vee\zeta_{n,\mu}(c_{3}))$, and so we consider only $\zeta_{n,\mu}(s)$ for $s<c_{3}$. To this end, using the fact that $\zeta_{n,\mu}$ are non-increasing, we have that
$\zeta_{n,\mu}(s)\leq\gamma/3\implies t\gtrsim\frac{c_{2}}{c_{1}}\zeta_{n,\mu}^{-1}(\gamma/3).$
@@
Before proving Theorems 3.5 and 3.6, we first prove two results concerning the projected quantiles and spacings. Let $\lambda_{t,q_{1},q_{2}}=\inf_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ z\in(\xi_{q_{1},u}-t,\xi_{q_{2},u}+t)\end{subarray}}f_{\nu_{u}}(z)$ and $\hat{\xi}_{q,u}=F^{-1}_{\hat{\nu}_{u}}(q)$.
@@
We now consider the event $E_{9}$. Recall that the set $\mathcal{N}_{\epsilon}$ is an $\epsilon$-net on $\mathbb{S}^{d-1}$ if for all $u\in\mathbb{S}^{d-1}$, there exists $v\in\mathcal{N}_{\epsilon}$ such that $\lVert u-v\rVert\leq\epsilon$. Letting $\mathcal{N}_{\epsilon}$ be an arbitrary $\epsilon$-net on $\mathbb{S}^{d-1}$, it is straightforward (see, e.g., [30, Lemma 4.4.1]) to show that for any $i,j\in[n]$,
@@
Furthermore, there exists $\mathcal{N}_{1/4}$ such that $|\mathcal{N}_{1/4}|\leq 9^{d}$ [30, see Corollary 4.2.13]. Applying these two facts, in conjunction with two applications of subadditivity of probability measures, yields